the total age between a father and son is 56 years.14 years later, the fathers age is twice of the son`s age. how old is the son?
Answer by richwmiller(17219)

The question is unclear.
Is the sum of the ages 56?
the sum is 56
f+s=56
f+14=2*(s+14)
f=42 s=14
check
56=2 *28
ok
or is the difference between the ages 56?
the difference is 56
f-s=56
f+14=2*(s+14)
f=98
s=42
check
98+14=112
42+14=56
112=2*56
ok